Two liquids A and B form an ideal solution. What is the vapour pressure of solution containing 2 moles of A and 3 moles of B at 300 K? (Vapour pressure of pure liquid A and B are 100 torr and 300 torr respectively) 

Grade:12

2 Answers

m aasif
6 years ago
P of solution = PureP of X . Molefraction of X + PureP of Y. Mole fractionof Y For the firstequation 500= Px .1/4 +Py . 3/4 Otherequation 600= Px.1/5 + Py . 2/5 Nowsolve both theequations and getthe answerDear Student,Please find belowthe solution to theasked query:Total pressure P= P X + P Xywhere = ​X =mole fraction of XX = molefraction of Y ​P =pure vapour pressure of XP = purevapour pressure ofYP =pressure of thetotal mixtureFirst condition = 500=Pox[11+3] + Poy[31+3] 500 = Pox4+3Poy4200Second condition 600 = Pox[11+4] +Poy[41+4]600 =Pox5+4Poy53000So we get twoequations as2000 = P + 3P3000 = P +4PSubtracting-1000 = -PP = 1000 mm HgP = -1000 mm HgHope thisinformation willclear your doubtsabout the topic
m aasif
6 years ago
P of solution = PureP of X . Molefraction of X + PureP of Y. Mole fractionof Y For the firstequation 500= Px .1/4 +Py . 3/4 Otherequation 600= Px.1/5 + Py . 2/5 Nowsolve both theequations and getthe answerDear Student,Please find belowthe solution to theasked:Total pressure P= P X + P Xywhere = ​X =mole fraction of XX = molefraction of Y ​P =pure vapour pressure of XP = purevapour pressure ofYP =pressure of thetotal mixtureFirst condition = 500=Pox[11+3] + Poy[31+3] 500 = Pox4+3Poy4200Second condition 600 = Pox[11+4] +Poy[41+4]600 =Pox5+4Poy53000So we get twoequations as2000 = P + 3P3000 = P +4PSubtracting-1000 = -PP = 1000 mm HgP = -1000 mm HgHope thisinformation willclear your doubtsabout the topic